# The Muckenhoupt condition

### v2

@@ 3.3. Bounds of the Hardy-Little maximal operator
Moreover, for a collection of cubes $\mathcal{P}$ we set $M^{\mathcal{P}}f:=\sup_{Q\in\mathcal{P}}T_{Q}f$. The generalized Muckenhoupt conditions are directly related to boundedness properties of $M$. Indeed, if $X$ is a Banach function space over $\mathbf{R}^{d}$ with the Fatou property, we have the following chain of implications:
$M:X$ $\to X,\,M:X^{\prime}\to X^{\prime}\Rightarrow X\in A_{\text{sparse}}\Rightarrow M:X\to X$
$\Rightarrow X\in A_{\text{strong}}\Rightarrow M:X\to X_{\text{weak}}\Rightarrow X\in A$
with
(3.7) $\begin{split}[X]_{A}&\leq\|M\|_{X\to X_{\text{weak}}}\lesssim_{d}[X]_{A_{\text{strong}}}\leq\|M\|_{X\to X}\\ &\lesssim_{d}[X]_{A_{\text{sparse}}}\lesssim\|M\|_{X\to X}\|M\|_{X^{\prime}\to X^{\prime}}.\end{split}$
More precisely, we have the following result.
@@
$[X]_{A}=\sup_{Q}\|T_{Q}\|_{X\to X_{\text{weak}}}\leq\|\sup_{Q}T_{Q}\|_{X\to X_{\text{weak}}}=\|M\|_{X\to X_{\text{weak}}},$
@@
For b, by the $3^{d}$-lattice theorem, it suffices to prove the result for $M^{\mathcal{D}}$ for a dyadic grid $\mathcal{D}$. Moreover, by the Fatou property of $X_{\text{weak}}$, it suffices to prove the result for $M^{\mathcal{F}}$ for a finite collection $\mathcal{F}\subseteq\mathcal{D}$. Let $\lambda>0$, $f\in X$, and pick a pairwise disjoint collection $\mathcal{P}\subseteq\mathcal{F}$ so that $\langle f\rangle_{1,P}>\lambda$ for all $P\in\mathcal{P}$ and
@@
Then
@@
$[X]_{A_{\text{sparse}}}\lesssim\|M\|_{X\to X}\|M\|_{X^{\prime}\to X^{\prime}}.$
@@
The assertion (3.10) then follows from the Fatou property of $X$.
@@
$[X]_{A}\leq\|M\|_{X\to X_{\text{weak}}}\lesssim_{d}[X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
@@ 3.4. $A_{p}$-regularity
To streamline our notation, we will call a collection of weights $B$ a weight class if there is an associated constant $[\cdot]_{B}$ with the property that $w\in B$ if and only if $[w]_{B}<\infty$.
@@
For a, let $g\in[(X^{p})^{\prime}]^{\frac{1}{p}}$ and $f\in X^{p}$ with $\|f\|_{X^{p}}=1$. Picking a weight $w\in A_{p}$ with $w\geq|f|^{\frac{1}{p}}$, $\|w\|_{X}\leq C_{1}$, $[w]_{p}\leq C_{2}$, we have
@@
For b, let $f\in X\cdot L^{p^{\prime}}(\mathbf{R}^{d})$, pick $0\leq h\in X$, $0\leq k\in L^{p^{\prime}}(\mathbf{R}^{d})$ such that $|f|\leq hk$, and pick $w\geq h$ as in the definition of $A_{p}$ regularity. Then, using Buckley’s bound [Buc93] and the fact that $[w^{-1}]_{p^{\prime}}=[w]_{p}$, we have
@@
[…]
Theorem 3.20 follows directly from a characterization proven by Rutsky in [Rut15]. We prove a sharp version of it.
@@
    $[(X^{r})^{\prime}]^{\frac{1}{r}}$ is $A_{r}$-regular.
@@
The main ingredient in the proof is a version of a classical result by Rubio de Francia [Rub86]. The following version is proven in [ALV19, Lemma 3.4] (see also [Lor16, Corollary 6.1.4]).
@@
Let $r\geq 1$ and let $X$ be an $r$-convex Banach function space over $\mathbf{R}^{d}$ with the Fatou property. Let $\Gamma$ be a family of (sub)linear operator on $X$. Then the following are equivalent:
@@
    $[(X^{r})^{\prime}]^{\frac{1}{r}}$ is $B$-regular with $C_{1}=2^{\frac{1}{r}}$, where
    $[w]_{B}:=\sup_{T\in\Gamma}\|T\|_{L^{r}_{w}(\mathbf{R}^{d})\to L^{r}_{w}(\mathbf{R}^{d})},\quad B:=\{w:[w]_{B}<\infty\}.$
@@
Instead of the Fatou property (which is assumed by Rubio de Francia [Rub86]), one can assume that $X$ is order-continuous (as is done, for example, in the versions of this result by Rutsky [Rut14, Rut19]). Actually, the only requirement for the result is that $X^{\prime}$ is norming for $X$.
@@
[…]
@@
[…]
$[w]_{2}\lesssim\|T\|_{L^{2}_{w}(\mathbf{R}^{d})\to L^{2}_{w}(\mathbf{R}^{d})}\lesssim\|\widetilde{T}\|_{X[\ell^{2}]\to X[\ell^{2}]}\lesssim\|T\|_{X\to X}.$
@@
$[X]_{A_{\text{sparse}}}\lesssim_{d}\|T\|_{X\to X}^{2}.$
@@
$[X]_{A_{\text{sparse}}}\lesssim_{d}C^{2}.$
@@ 4. Duality of the Hardy-Littlewood maximal operator
The general duality question (Q) is posed for Banach function spaces rather than the more general quasi-Banach function spaces. It is worth noting that it is not true that if $M$ is bounded on a quasi-Banach function space $X$, then $X$ can be renormed to a Banach function space. For example, the Lorentz space $X=L^{p,q}(\mathbf{R}^{d})$ for $p\in(1,\infty)$ and $q\in(0,1)$ is a quasi-Banach function space on which $M$ is bounded that is not a Banach function space. However, we claim that solving (Q) as posed is sufficient to also conclude the result for any quasi-Banach function space $X$. Indeed, if $M$ is bounded on $X$, then $X^{\prime}$ is a Banach function space, and thus $X^{\prime\prime}$ is also a Banach function space. It is actually possible to show that then $M$ is also bounded on the Banach function space $X^{\prime\prime}$. Since $X^{\prime\prime\prime}=X^{\prime}$, we can now simply replace $X$ by $X^{\prime\prime}$ in (Q) to obtain the desired conclusion.
This duality result was originally shown in [Nie23, Proposition 2.27], and we give a brief proof based on the linearization $\mathcal{M}$ here.
@@
[…]
@@
for all $p>1$ and all $f,g\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$. This can be used to establish that $M:(X^{\prime})^{\theta}\to(X^{\prime})^{\theta}$ for any $10<\theta<$ 1. As a matter of fact, we have the following general duality results.
@@
    $M:(X^{\prime})^{\theta}\to(X^{\prime})^{\theta}$ for all $10<\theta<$ 1 with
@@
    $M:[(X^{\prime})^{\theta}]^{\prime}\to[(X^{\prime})^{\theta}]^{\prime}$ for all $10<\theta<$ 1 with
@@
    $(X^{\prime})^{\theta}\in A_{\text{sparse}}$ for all $10<\theta<$ 1 with
    $[(X^{\prime})^{\theta}]_{A_{\text{sparse}}}\lesssim_{d}\frac{1}{\theta}\frac{1}{1-\theta}\|M\|_{X\to X}^{1+\theta}.$
@@
For b, note that by (2.1) we have $[(X^{\prime})^{\theta}]^{\prime}=X^{\frac{1}{p}}\cdot L^{p^{\prime}}(\mathbf{R}^{d})$. Thus, the result follows from Proposition 3.19 applied to the space $X^{\frac{1}{p}}$, using the observation that since $X$ is $A_{1}$-regular with constant $C_{2}\lesssim\|M\|_{X\to X}$ and for any $w\in A_{1}$ we have
$[w^{\frac{1}{p}}]^{p}_{p}\leq[w]_{1}$
@@
To statement c follows from Theorem 3.14e applied to the space $(X^{\prime})^{\theta}$, combined with a and b. The assertion follows. ∎
We point out that a and b in Theorem 4.2 imply that if $M:X\to X$, then, for any $10<\theta<$ 1, the space $Y:=(X^{\prime})^{\theta}$ satisfies
@@
This observation results in the following characterizations for boundedness of $M$ on a space and its Köthe dual, which were originally proven in [LN24b]:
@@
    there is a $C>0$ such that for all $f\in X$ satisfying |$|\{x\in\mathbf{R}^{d}:|f(x)|>\lambda\}|<\infty$ for all $\lambda>0$ we have
@@
    $M:(X^{\theta})^{\prime}\to(X^{\theta})^{\prime}$ for some $10<\theta<$ 1;
@@
[…]
@@
that satisfies $\phi_{X^{p}}=(\phi_{X})^{p}$ for any $p>0$. In particular, this means that $\Leftrightarrow$ $\phi_{X}$ is unbounded $\Leftrightarrow$ $\phi_{X^{\theta}}$ is unbounded. Since $M:X\to X$ also implies $M:X^{\theta}\to X^{\theta}$, this last assertion is equivalent to iv, as desired.
@@ 5.1. Unweighted variable Lebesgue spaces
    $L^{p(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ with $[X]_{A_{\text{strong}}}\lesssim_{d,C_{1},C_{2}}1$;
@@
The first result is proven in [DHHR11, Theorem 4.4.8]. Note that by Theorem 3.14b this also implies that
@@
The second result was proven in [CFN03]. See also [DHHR11, Theorem 4.3.8], and the paragraph that follows it for the history of this problem.
@@ 5.2. Weighted variable Lebesgue spaces
$[L^{p(\cdot)}_{w}(\mathbf{R}^{d})]_{A}=\sup_{Q}\|w\operatorname{\mathbf{1}}_{Q}\|_{L^{p(\cdot)}(\mathbf{R}^{d})}\|w^{-1}\operatorname{\mathbf{1}}_{Q}\|_{L^{p^{\prime}(\cdot)}(\mathbf{R}^{d})},$
@@
The equivalence of the last statement with iii is proven in [CDH11, Theorem 1.3]. For the equivalence of i-iii, we first note that by [DHHR11, Theorem 7.3.22], global $\log$-Hölder continuity implies that $L^{p(\cdot)}(\mathbf{R}^{d})\in\mathcal{G}$. Following the observation of [CDH11, Lemma 2.1] (which notes that if $X\in\mathcal{G}$ holds in the unweighted space, then it automatically also hold in the weighted space by applying the unweighted estimate to $fw$, $gw^{-1}$ instead of $f$, $g$), we conclude that for any weight $w$ we have $L^{p(\cdot)}_{w}(\mathbf{R}^{d})\in\mathcal{G}$ with
$[L^{p(\cdot)}_{w}(\mathbf{R}^{d})]_{\mathcal{G}}=[L^{p(\cdot)}(\mathbf{R}^{d})]_{\mathcal{G}}.$

### v3

@@ 3.3. Bounds of the Hardy-Little maximal operator
Moreover, for a collection of cubes $\mathcal{P}$ we set $M^{\mathcal{P}}f:=\sup_{Q\in\mathcal{P}}T_{Q}f$. The generalized Muckenhoupt conditions are directly related to boundedness properties of $M$.
@@
$X]_{A}=\sup_{Q}\|T_{Q}\|_{X\to X_{\text{weak}}}\leq\|\sup_{Q}T_{Q}\|_{X\to X_{\text{weak}}}=\|M\|_{X\to X_{\text{weak}}},$
@@
For b, by the $3^{d}$-lattice theorem, it suffices to prove the result for $M^{\mathcal{D}}$ for a dyadic grid $\mathcal{D}$. Moreover, by the Fatou property of $X_{\text{weak}}$, it suffices to prove the result for $M^{\mathcal{F}}$ for a finite collection $\mathcal{F}\subseteq\mathcal{D}$. Let $\lambda>0$, $f\in X$, and let $\mathcal{P}$ denote the collection of maximal cubes $P\in\mathcal{F}$ satisfying $\langle f\rangle_{1,P}>\lambda$, such that we have the Calderón-Zygmund decomposition
@@
As $\mathcal{P}$ is pairwise disjoint, we have $\operatorname{\mathbf{1}}_{\{x\in\mathbf{R}^{d}:M^{\mathcal{F}}f(x)>\lambda\}}=\sum_{P\in\mathcal{P}}\operatorname{\mathbf{1}}_{P}$, which implies that
@@
$X]_{A_{\text{sparse}}}\lesssim\|M\|_{X\to X}\|M\|_{X^{\prime}\to X^{\prime}}.$
@@
The assertion (3.9) then follows from the Fatou property of $X$. In the case $X=L^{1}(\mathbf{R}^{d})$, (3.9) is similar to a more classical estimate (with $M^{\mathcal{D}}f$ replaced by $2f$ in the upper level set on the right-hand side) which is part of an equivalence that can be found, e.g., in [GR85, p. 144, Theorem 2.1]. This opens up the question of whether this same equivalence also holds for general $X$.
@@
$X]_{A}\leq\|M\|_{X\to X_{\text{weak}}}\lesssim_{d}[X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
@@ 3.4. $A_{p}$-regularity
To streamline our notation, we will call a collection of weights $B$ a weight class if there is an associated constant $\cdot]_{B}$ with the property that $w\in B$ if and only if $w]_{B}<\infty$.
@@
For a, let $g\in[(X^{p})^{\prime}]^{\frac{1}{p}}$ and $f\in X^{p}$ with $\|f\|_{X^{p}}=1$. Picking a weight $w\in A_{p}$ with $w\geq|f|^{\frac{1}{p}}$, $\|w\|_{X}\leq C_{1}$, $w]_{p}\leq C_{2}$, we have
@@
For b, let $f\in X\cdot L^{p^{\prime}}(\mathbf{R}^{d})$, pick $0\leq h\in X$, $0\leq k\in L^{p^{\prime}}(\mathbf{R}^{d})$ such that $|f|\leq hk$, and pick $w\geq h$ as in the definition of $A_{p}$ regularity. Then, using Buckley’s bound [Buc93] and the fact that $w^{-1}]_{p^{\prime}}=[w]_{p}$, we have
@@
Using Theorem 3.26 below, we point out that the implication $i\Rightarrowi$ is true for $r=2$ without any convexity assumption on $X$.
Theorem 3.21 follows directly from a characterization proved by Rutsky in [Rut15]. We prove a sharp version of it.
@@
    $(X^{r})^{\prime}]^{\frac{1}{r}}$ is $A_{r}$-regular.
@@
The main ingredient in the proof is a version of a classical result by Rubio de Francia [Rub86]. The following version is proved in [ALV19, Lemma 3.4] (see also [Lor16, Corollary 6.1.4]).
@@
Let $r\geq 1$ and let $X$ be an $r$-convex Banach function space over $\mathbf{R}^{d}$ with the Fatou property. Let $\Gamma$ be a family of (sub)linear operators on $X$. Then the following are equivalent:
@@
    $(X^{r})^{\prime}]^{\frac{1}{r}}$ is $B$-regular with $C_{1}=2^{\frac{1}{r}}$, where
    $w]_{B}:=\sup_{T\in\Gamma}\|T\|_{L^{r}_{w}(\mathbf{R}^{d})\to L^{r}_{w}(\mathbf{R}^{d})},\quad B:=\{w:[w]_{B}<\infty\}.$
@@
Instead of the Fatou property (which is assumed by Rubio de Francia [Rub86]), one can assume that $X$ is order-continuous (as is done, for example, in the versions of this result by Rutsky [Rut14, Rut19]). Actually, the only requirement for the result is that $X^{\prime}$ is norming for $X$, i.e., for all $f\in X$ we have $\|f\|_{X}=\sup_{\|g\|_{X^{\prime}}=1}\int_{\mathbf{R}^{d}}\!|fg|\,\mathrm{d}x$.
@@
The implication $i\Rightarrowiii$ follows from Proposition 3.20, where the Fatou property of $X$ is used so that $Y=[(X^{r})^{\prime}]^{\frac{1}{r}}$ satisfies $(Y^{r})^{\prime}]^{\frac{1}{r}}=[(X^{r})^{\prime\prime}]^{\frac{1}{r}}=X$ by the Lorentz-Luxemburg theorem.
@@
First, we show that $\Rightarrowvi\Rightarrow i$v, proving the equivalence of vi with i-iv.
@@
$w]_{2}\lesssim\|T\|_{L^{2}_{w}(\mathbf{R}^{d})\to L^{2}_{w}(\mathbf{R}^{d})}\lesssim\|\widetilde{T}\|_{X[\ell^{2}]\to X[\ell^{2}]}\lesssim\|T\|_{X\to X}.$
@@
$X]_{A_{\text{sparse}}}\lesssim_{d}\|T\|_{X\to X}^{2}.$
@@
$X]_{A_{\text{sparse}}}\lesssim_{d}C^{2}.$
@@ 4. Duality of the Hardy-Littlewood maximal operator
The general duality question (Q) is posed for Banach function spaces rather than the more general quasi-Banach function spaces. It is worth noting that it is not true that if $M$ is bounded on a quasi-Banach function space $X$, then $X$ can be renormed to a Banach function space. For example, the Lorentz space $X=L^{p,q}(\mathbf{R}^{d})$ for $p\in(1,\infty)$ and $q\in(0,1)$ is a quasi-Banach function space on which $M$ is bounded that is not a Banach function space. However, we claim that solving (Q) as posed is sufficient to also conclude the result for any quasi-Banach function space $X$.
Indeed, suppose $X$ is a quasi-Banach function space on which $M$ is bounded. Then, by a duality result originally shown in [Nie23, Proposition 2.27], $M$ is also bounded on $X^{\prime\prime}$, which is a Banach function space. Since $X^{\prime\prime\prime}=X^{\prime}$, we need only answer (Q) for $X^{\prime\prime}$ instead of $X$, as claimed.
We give a brief proof of this duality result based on the linearization $\mathcal{M}$ here.
@@
But, with the same proof as the one of $i\Rightarrowiv i$n Proposition 2.3, this follows from $M:X\to X$ with
@@
for all $p>1$ and all $f,g\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$. This can be used to establish that $M:(X^{\prime})^{\theta}\to(X^{\prime})^{\theta}$ for any $0<\theta<1$. As a matter of fact, we have the following general duality results.
@@
    $M:(X^{\prime})^{\theta}\to(X^{\prime})^{\theta}$ for all $0<\theta<1$ with
@@
    $M:[(X^{\prime})^{\theta}]^{\prime}\to[(X^{\prime})^{\theta}]^{\prime}$ for all $0<\theta<1$ with
@@
    $(X^{\prime})^{\theta}\in A_{\text{sparse}}$ for all $0<\theta<1$ with
    $(X^{\prime})^{\theta}]_{A_{\text{sparse}}}\lesssim_{d}\frac{1}{\theta}\frac{1}{1-\theta}\|M\|_{X\to X}^{1+\theta}.$
@@
For b, note that by (2.1) we have $(X^{\prime})^{\theta}]^{\prime}=X^{\frac{1}{p}}\cdot L^{p^{\prime}}(\mathbf{R}^{d})$. Thus, the result follows from Proposition 3.20 applied to the space $X^{\frac{1}{p}}$, using the observation that since $X$ is $A_{1}$-regular with constant $C_{2}\lesssim\|M\|_{X\to X}$ and for any $w\in A_{1}$ we have
$w^{\frac{1}{p}}]^{p}_{p}\leq[w]_{1}$
@@
Statement c follows from Theorem 3.15e applied to the space $(X^{\prime})^{\theta}$, combined with a and b. The assertion follows. ∎
We point out that a and b in Theorem 4.2 imply that if $M:X\to X$, then, for any $0<\theta<1$, the space $Y:=(X^{\prime})^{\theta}$ satisfies
@@
This observation results in the following characterizations for boundedness of $M$ on a space and its Köthe dual, which were originally proved in [LN24b]:
@@
    there is a $C>0$ such that for all $f\in X$ satisfying $|\{x\in\mathbf{R}^{d}:|f(x)|>\lambda\}|<\infty$ for all $\lambda>0$ we have
@@
    $M:(X^{\theta})^{\prime}\to(X^{\theta})^{\prime}$ for some $0<\theta<1$;
@@
The characterization $\Leftrightarrowiii i$s contained in the work of Lerner [Ler10, Corollary 4.3]. Perhaps interestingly, one can also prove the implication $i\Rightarrowiii$ directly by using the weighted Fefferman-Stein inequality
@@
[…]
@@ 5.1. Unweighted variable Lebesgue spaces
    $L^{p(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ with $X]_{A_{\text{strong}}}\lesssim_{d,C_{1},C_{2}}1$;
@@
The first result is proved in [DHHR11, Theorem 4.4.8]. Note that by Theorem 3.15b this also implies that
@@
The second result was proved in [CFN03]. See also [DHHR11, Theorem 4.3.8], and the paragraph that follows it for the history of this problem.
@@ 5.2. Weighted variable Lebesgue spaces
$L^{p(\cdot)}_{w}(\mathbf{R}^{d})]_{A}=\sup_{Q}\|w\operatorname{\mathbf{1}}_{Q}\|_{L^{p(\cdot)}(\mathbf{R}^{d})}\|w^{-1}\operatorname{\mathbf{1}}_{Q}\|_{L^{p^{\prime}(\cdot)}(\mathbf{R}^{d})},$
@@
The equivalence of the last statement with iii is proved in [CDH11, Theorem 1.3]. For the equivalence of i-iii, we first note that by [DHHR11, Theorem 7.3.22], global $\log$-Hölder continuity implies that $L^{p(\cdot)}(\mathbf{R}^{d})\in\mathcal{G}$. Following the observation of [CDH11, Lemma 2.1] (which notes that if $X\in\mathcal{G}$ holds in the unweighted space, then it automatically also hold in the weighted space by applying the unweighted estimate to $fw$, $gw^{-1}$ instead of $f$, $g$), we conclude that for any weight $w$ we have $L^{p(\cdot)}_{w}(\mathbf{R}^{d})\in\mathcal{G}$ with
$L^{p(\cdot)}_{w}(\mathbf{R}^{d})]_{\mathcal{G}}=[L^{p(\cdot)}(\mathbf{R}^{d})]_{\mathcal{G}}.$